OpenStudy (anonymous):

solve each equation and show all work 1. 5x+7=3x+21 2. 3x-2(5-x)=3(x-10)+3 3. 5(x+1)=3(2x+3)+5

OpenStudy (anonymous):

Line up everything that matches. like variables.

OpenStudy (anonymous):

So for #1 5x+7=3x+21 so what would you do?

OpenStudy (anonymous):

subtract 3x from both sides?

OpenStudy (gabbyalicorn):

Correct!

OpenStudy (anonymous):

Correct you want to get rid of one of the x's. So you would get 2x+7=21 What do you do next?

OpenStudy (anonymous):

subtract 7 from both sides

OpenStudy (anonymous):

Subtract 7 from both sides you get 14 so it would be 2x=14 divide both sides by 2 and you get x=7! Now do the rest :)

OpenStudy (anonymous):

Hint: parentheses first always remember that.

OpenStudy (anonymous):

how would i start the ones in the parentheses

OpenStudy (anonymous):

3x-2(5-x)=3(x-10)+3 Distribute. -2times 5 and -2 times x you would also do 3 times x and 3 times -10 your equation should be 3x-10-2x=3x-30+3

OpenStudy (anonymous):

so do you get rid of the 3x first

OpenStudy (anonymous):

No the -2x so add 2x to the first 3x you get 5x-10=3x-30+3 then subtract 3x so you get 2x-10=-30+3 Subtract 3 you get 2x-13+-30 add 30 you get 2x=20 divide by 2 you get x=10 :)

OpenStudy (anonymous):

#3 answer is -9

OpenStudy (anonymous):

how did you get -9

ILovePuppiesLol (ilovepuppieslol):

which one are you doing rn?

ILovePuppiesLol (ilovepuppieslol):

@kiki345

OpenStudy (anonymous):

# 3

ILovePuppiesLol (ilovepuppieslol):

alright

ILovePuppiesLol (ilovepuppieslol):

\( 5(x+1)=3(2x+3)+5\) open the brackets \(5(x)+5(1)=3(2x)+3(3)+5\) \(5x+5=6x+9+5\) \(-1x=9\) \(x=-9\)

OpenStudy (anonymous):

thank you @ILovePuppiesLol

OpenStudy (anonymous):

I can

OpenStudy (anonymous):

-2(6-2x)=4(-3+x)

OpenStudy (anonymous):

−2(6−2x)=4(−3+x) Step 1: Simplify both sides of the equation. −2(6−2x)=4(−3+x) Simplify: (Show steps) 4x−12=4x−12 Step 2: Subtract 4x from both sides. 4x−12−4x=4x−12−4x −12=−12 Step 3: Add 12 to both sides. −12+12=−12+12 0=0

OpenStudy (anonymous):

thank you and i have one more 5-1(2x+3)=-2(4+x)

OpenStudy (anonymous):

5−1(2x+3)=−2(4+x) Step 1: Simplify both sides of the equation. 5−1(2x+3)=−2(4+x) Simplify: (Show steps) −2x+2=−2x−8 Step 2: Add 2x to both sides. −2x+2+2x=−2x−8+2x 2=−8 Step 3: Subtract 2 from both sides. 2−2=−8−2 0=−10

OpenStudy (anonymous):

thanks